# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ V Toward high-performance hybrid quantum–classical computing
Quantum computing has generated considerable interest in the high-performance computing (HPC) community as a promising extension beyond exascale systems. As accelerators within HPC infrastructures, quantum computers could perform specialized tasks rather than replacing classical computers as general-purpose systems. To reach utility-scale quantum computing, seamless integration with existing heterogeneous HPC infrastructures and the development of a full hybrid quantum–classical stack are essential.
@@
Integrating quantum computing with high-performance computing (QC-HPC) presents several challenges. On the hardware and system design fronts, key differences between quantum and classical components include physical scale, reliability, control electronics, communication bandwidth, and operational time scales. Algorithmically, the challenges involve memory access, data sharing and movement, and efficient information extraction. Some quantum algorithms lack clearly defined kernels to be off-loaded to QPUs. For certain hybrid quantum–classical algorithms, the data movement overhead associated with offloading portions to a quantum device could diminish or erase performance gains the quantum algorithm could in theory provide. This is especially true for variational algorithms, where the quantum kernel is executed multiple times in tight interaction with a classical program. These challenges must be factored into the practical design and implementation of a hybrid quantum-HPC system. Physically co-locating classical and quantum computing resources within the same hardware node might be necessary when classical and quantum components need to exchange data with tight latency or require frequent synchronizations. To enable low-latency high-bandwidth communication, QPUs should be tightly interconnected with multiple CPUs (cores) and other accelerators such as GPUs and FPGAs, all sharing the same system resources such as memory, cache, and high-speed interconnects.
@@
Beyond physical integration, it is necessary to ensure the hybrid quantum-HPC system is easily programmable for the end user. Considering QPUs as accelerators and aiming for minimal changes to overall HPC program structure can mitigate the risks of complicated system development with specialized hardware. A natural solution is to integrate tools that program, compile, and execute quantum circuits into current classical HPC programming environments. Existing infrastructure for HPC (e.g. data and user management, process scheduling, control and networking) can then be leveraged for future quantum-HPC systems. For many end users, access at the HPC programming environment level will be familiar and best. More advanced users, however, may want access to the underlying physical hardware and cyber-physical control system. Different levels of abstraction in the quantum-HPC software portfolio should be harnessed for the different needs of the end users.
@@
Figure 1 illustrates the architecture diagram for a comprehensive HPC software portfolio with extensions toward a full quantum-HPC stack. The HPE Cray Programming Environment (CPE) is a mature HPC programming system that provides software development toolchains supporting a full range of heterogenous HPC platforms, hardware architectures, and processors. CPE provides support for multiple programming environments including HPE Cray, AMD, Intel, Nvidia, and GNU with compiler interoperability. Building on top of CPE can significantly reduce development efforts and enable rapid experimentation with different quantum SDKs (e.g., CUDA-QNvidia. , QiskitIBM. , CirqGoogle. , PennylaneXanadu. , and ClassiqClassiq Technologies Ltd. ) on available quantum and quantum-inspired accelerators as well as simulators. We can identify and target modular software capable of adapting to emerging and increasingly powerful QPU technologies, while at the same time leveraging existing NISQ QPUs as well as CPU/GPU cores for high-performance simulation. The following section describes these extensions to the HPC programming environment in further detail.
@@ V.1 HPC programming environment extensions
Figure 30: Schematic of a hybrid quantum-HPC application development workflow utilizing the CPE quantum interface library. HPC applications in Python or compiled languages (e.g. C/C++) can call different quantum SDKs (via quantum interface library) and still link with other libraries, e.g. Cray Science and Math Library (CSML). Circuit synthesis is then available from a variety of quantum SDKs—CUDA-Q, Qiskit, Pennylane, Classiq, etc. Synthesized quantum circuits can be subsequently executed on various supported quantum hardware and simulators hosted remotely on cloud or on-premise. Simulation results are sent back as return values to the HPC applications.
@@
With diverse quantum hardware including superconducting qubits, trapped ions, neutral atoms, and photonic qubits, various quantum software packages focus on different aspects of quantum computing. This may include circuit synthesis or optimizing complex design processes including qubit allocation, auxiliary qubit reuse, error mitigation, or quantum error correction. These quantum software packages are developed in different programming models and have parallelism models supporting different backends. For example, CUDA-Q supports both task-based and distributed parallel circuit execution models and provides multi-GPU, multi-node state vector and tensor network simulation backends on Nvidia GPUsNVIDIA. . Pennylane Lightning, as another example, provides state vector simulators that can be executed on both AMD and Nvidia GPUsAsadi et al. (2024).
@@
To enable seamless invocation of quantum kernels from different quantum SDKs within HPC applications developed in C/C++/Fortran, a quantum API-CPE Quantum Interface Library (CPE-QIL) is implemented in C, which can be seamlessly interfaced with Fortran applications using the iso_c_binding module. Application developers can use high-level, portable invocation of quantum algorithm libraries from a variety of third party SDKs within a classical HPC application in their programming language of choice. Figure 30 provides a schematic for a hybrid quantum-HPC development workflow within CPE with the quantum interface library.
@@
Circuit synthesis and execution time in quantum computing can vary significantly based on the complexity of the algorithm and the number of qubits involved. Different SDKs offer various levels of optimization and distinct approaches to handling quantum gates, scheduling, and resource allocation, all of which impact the time and efficiency of circuit synthesis and execution. Each quantum SDK (e.g. Qiskit, CUDA-Q, etc.) has its native gate set optimized for the underlying hardware. The choice of gate set impacts synthesis complexity since some SDKs might require multiple gates to synthesize specific operations efficiently, while others may directly support them. Decomposition of high-level operations into hardware-native gate sets is usually required. For example, synthesizing a rotation or controlled operation might require multiple CNOT gates and rotations, affecting both gate count and execution time. Many SDKs provide optimization levels to minimize circuit depth, gate count, or overall execution time. These optimizations directly influence both the synthesis time (by making trade-offs for synthesis overhead) and execution time on supported simulators and hardware. An example of time ranges is given in Section VI.1, where circuit synthesis for Trotterization is found to take on the order of milliseconds.
@@
Quantum interface library with API extensions: The quantum interface library extension within CPE has the following advantages: 1) the ability to support and interact with a range of quantum SDKs and backends from different vendors through a standardized interface, 2) data and functionality of other software systems are available while implementation details are abstracted, facilitating efficient and secure application development, 3) support for compiled languages commonly used in massively parallel HPC applications which could allow large-scale system modeling and computation with large data sets, and 4) direct utilization of the existing HPE Cray MPI, which offers "GPU aware" MPI support and heterogeneous workload and resource managers such as Slurm and PBS (Portable Batch System) Slurm. ; OpenPBS. .
@@
The CPE quantum interface library has been utilized to deliver two hybrid quantum-HPC applications in both Python and C/C++ with quantum kernels for circuit synthesis and execution provided by Classiq’s Python based quantum SDK Classiq Technologies Ltd. . One example consists of solving linear systems of equations with the Harrow-Hassidim-Lloyd (HHL) algorithmJose Morrell Jr et al. (2021). The solution was compared with the CPE BLAS library and showed less than 2% deviation. Another example considers the quantum approximate optimization algorithm (QAOA) for solving the MaxCut problem of partitioning a large graph into smaller sub-graphs, which can each be represented on current quantum devices. The workflow is well suited for a hybrid classical-quantum execution on supercomputers, where various sub-problems can also be solved classically if there is an advantage. Initial investigations have been conducted by HPE and Classiq and published at IPDPS24 Esposito and Danzig (2024). For a simplified problem, this hybrid workload was demonstrated at ISC24 using a 20 qubit IQM quantum device accessed from the LUMI supercomputer in Finland. Understanding the latency implications in accessing a remote quantum device was one of the main goals of this investigation; the communication overhead was typically found to be on the order of seconds. This latency could be reduced for tightly integrated machines.
@@
Figure 31 presents a detailed illustration of hybrid quantum-HPC application development and execution within CPE with the quantum interface library and HPC workload management. Within CPE, users can automatically utilize debugging and profiling tools as well as math and communication libraries with chosen compilers. These compilers (for Fortran, C, and C++) are designed to extract maximum performance from a variety of architectures like ARM and x86-64 and devices like AMD and Nvidia GPUs. In addition, users have access to the HPE Cray Message Passing Interface (MPI), a highly scalable implementation for collective communications. Applications developed in C/C++/Fortran are compiled and linked within CPE, potentially including other libraries such as the Cray Science and Math Libraries (CSML) if the HPC application requires it. The quantum interface library is linked as a shared library during the application build process. Hybrid executables can be submitted and scheduled for parallel execution by a workload manager such as Slurm or PBS. At runtime, the interface library routes the quantum API calls from the application to the respective vendor-specific quantum SDKs with appropriate data handling (e.g. parameters and return values). When a quantum API call provided by the interface library is invoked by the application at runtime:
@@
  • The quantum circuit is executed (either on a quantum device or simulator) and results are converted and passed back to the application as return values of the quantum API
@@
Figure 31: Development and execution of hybrid quantum-HPC applications in C/C++/Fortran through the Quantum Interface Library within the HPE Cray Programming Environment (CPE). A hybrid executable is dispatched to a workload manager (e.g. Slurm or PBS) for parallel execution. Remote visualization through web applications can be enabled with a different protocol and configurations for secure connection to clusters.
@@
Quantum compiler and runtime extensions: With increasing qubit counts, higher gate fidelity, and improved coherence times and scalability, compilation bottlenecks and latency between classical and quantum components become more apparent. When scaling to large circuit sizes with $\sim$100 or more qubits, declarative frameworks struggle to handle compilation bottlenecks and latency between classical and quantum components in the program, even for NISQ systems. Leveraging classical compilation tool chains such as Clang/LLVM is crucial for developing large-scale hybrid quantum-HPC workloads, and research efforts are moving in this direction Litteken et al. (2020); Chong et al. (2017).
@@
To enable performant language-level support for quantum constructs and to address bottlenecks in compile-time with increasing circuit size, we have on-going efforts to build extensions on top of classical compilation tool chains. Given the diversity of pulse-level quantum instructions by different quantum hardware vendors, the quantum assembly language OpenQASM and LLVM IR with its extension to Quantum Intermediate Representation (QIR) Microsoft. are adopted as a middle ground for comparability with different quantum hardware and software. We leverage codegen modules to emit machine-specific native code for target architectures based on the LLVM IR/QIR produced by different quantum software front ends such as CUDA-Q. CUDA-Q provides the NVQ++ compiler for quantum kernels lowering to QIR eventually, as well as a standard library of quantum algorithmic primitives with upcoming support for quantum error correction primitives. Generated code, when linked with appropriate quantum vendor-provided runtime libraries, can be executed on target quantum devices or simulator backends in a quantum-backend-retargetable manner.
@@
Dynamic code generation techniques are adopted to generate code at runtime based on specific characteristics of the target quantum devices. In doing so, the compiler can optimize the execution of quantum programs for different hardware generations or architectures. This low-level integration at the IR level will ensure hardware support, software compatibility, and optimal circuit compilation and execution performance for large-scale quantum-HPC workload development.
@@
Multi-QPU workload distribution hypervisor: The final extension to the HPC programming environment aims to tackle the problem of quantum workload distribution. For quantum computing to operate at scale, it will be necessary to efficiently parallelize over many QPUs, possibly of different hardware types, integrating them as coprocessors within an HPC framework. We enable this integration by developing a novel adaptive circuit knitting strategy serving as a hypervisor for classical and quantum communication between distributed classical and quantum compute nodes. Unlike traditional circuit knitting, this strategy uses machine learning at multiple levels to learn a quantum circuit decomposition capturing maximal quantum entanglement while enabling high-performance communication at scale. By integrating scalable message passing techniques within an adaptive adaptive circuit knitting approach, we could efficiently learn how to perform distributed quantum simulation or distributed quantum machine learning over quantum data generated by quantum processors themselves. We introduce this approach in the following section.
V.2 High-performance quantum workload distribution
Existing quantum processors have relatively few qubits with low gate fidelity. With the current state of technology, it is highly unlikely for NISQ devices to be scaled to tackle utility-scale problems. While proof-of-principle demonstrations of logical qubits are just starting to appear, even upcoming error-corrected quantum computers will be very small with respect to the number of logical qubits required for utility, according to all industrial roadmaps. Recent resource estimates for FTQC indicates that the required number of physical qubits are about 0.5M-2M for quantum dynamics, 1M-6M for quantum chemistry, and 6M-30M for integer factoring Beverland et al. (2022). At the same time, the largest quantum processors to date are on the order of hundreds of qubits, and even most optimistic roadmaps do not anticipate more than 100k qubits at a single QPU level. Consequently, to reach utility-scale quantum computing, efficiently distributing computation across multiple QPUs will be required.
Efficiently partitioning quantum systems has a rich history in quantum science White (1993); Meyer et al. (1990); Manthe (2008). In recent years, circuit knitting has emerged as a promising method to partition quantum circuits, the primary goal being to enable simulating large circuits on NISQ devices available today Peng et al. (2020); Piveteau and Sutter (2023). In circuit knitting, a measured observable is reconstructed by sampling sub-circuits of the original circuit multiple times. This partitioning was introduced as an error mitigation mechanism by using a quasi-probability decomposition to mimic the output of a large noiseless quantum circuit by a number of smaller noisy quantum circuits Temme et al. (2017); Piveteau et al. (2022). However, this reconstruction comes at an exponential cost in the number of samples that depends on the identity and number of gates that have been cut out Piveteau and Sutter (2023). While recent efforts have focused on reducing this exponential overhead Tang et al. (2021); Tang and Martonosi (2022); Basu et al. (2024), further work is necessary to demonstrate the practical advantage of circuit knitting.

### v2

@@ V Toward high-performance hybrid quantum–classical computing
Quantum computing has generated considerable interest in the high-performance computing (HPC) community as a promising extension beyond exascale systems. As accelerators within HPC infrastructures, quantum computers could perform specialized tasks rather than replacing classical computers outright as general-purpose systems. To reach utility-scale quantum computing, seamless integration with existing heterogeneous HPC infrastructures and the development of a full hybrid quantum–classical stack are essential.
@@
Integrating quantum computing with high-performance computing (quantum–HPC) presents several challenges. On the hardware and system design fronts, key differences between quantum and classical components include physical scale, reliability, control electronics, communication bandwidth, and operational time scales. Algorithmically, the challenges involve memory access, data sharing and movement, and efficient information extraction. Some quantum algorithms lack clearly defined kernels to be off-loaded to QPUs. For certain hybrid quantum–classical algorithms, the data movement overhead associated with offloading portions to a quantum device could diminish or erase performance gains the quantum algorithm could in theory provide. This is especially true for variational algorithms, where the quantum kernel is executed multiple times in tight interaction with a classical program. These challenges must be factored into the practical design and implementation of a hybrid quantum–HPC system. Physically co-locating classical and quantum computing resources within the same hardware node might be necessary when classical and quantum components need to exchange data with tight latency or require frequent synchronizations. To enable low-latency high-bandwidth communication, QPUs should be tightly interconnected with multiple CPUs (cores) and other accelerators such as GPUs and FPGAs, all sharing the same system resources such as memory, cache, and high-speed interconnects.
@@
Beyond physical integration, it is necessary to ensure the hybrid quantum–HPC system is easily programmable for the end user. Considering QPUs as accelerators and aiming for minimal changes to overall HPC program structure can mitigate the risks of complicated system development with specialized hardware. A natural solution is to integrate tools that program, compile, and execute quantum circuits into current classical HPC programming environments. Existing infrastructure for HPC (e.g., data and user management, process scheduling, control and networking) can then be leveraged for future quantum–HPC systems. For many end users, access at the HPC programming environment level will be familiar and best. More advanced users, however, may want access to the underlying physical hardware and cyber-physical control system. Different levels of abstraction in the quantum–HPC software portfolio should be harnessed for the different needs of the end users.
@@
Figure 2 illustrates the architecture diagram for a comprehensive HPC software portfolio with extensions toward a full quantum–HPC stack. The HPE Cray Programming Environment (CPE) is a mature HPC programming system that provides software development toolchains supporting a full range of heterogenous HPC platforms, hardware architectures, and processors. CPE provides support for multiple programming environments including HPE Cray, AMD, Intel, Nvidia, and GNU with compiler interoperability. Building on top of CPE can significantly reduce development efforts and enable rapid experimentation with different quantum SDKs (e.g., CUDA-Q NVIDIA. , Qiskit IBM. , Cirq Google. , Pennylane Xanadu. , and Classiq Classiq Technologies Ltd. ) on available quantum and quantum-inspired accelerators as well as simulators. We can identify and target modular software capable of adapting to emerging and increasingly powerful QPU technologies, while at the same time leveraging existing NISQ QPUs as well as CPU/GPU cores for high-performance simulation. The following section describes these extensions to the HPC programming environment in further detail.
@@ V.1 HPC programming environment extensions
Figure 32: Schematic of a hybrid quantum–HPC application development workflow utilizing the CPE quantum interface library. HPC applications in Python or compiled languages (e.g. C/C++) can call different quantum SDKs (via quantum interface library) and still link with other libraries, e.g., Cray Science and Math Library (CSML). Circuit synthesis is then available from a variety of quantum SDKs—CUDA-Q, Qiskit, Pennylane, Classiq, etc. Synthesized quantum circuits can be subsequently executed on various supported quantum hardware and simulators hosted remotely on cloud or on-premise. Simulation results are sent back as return values to the HPC applications.
@@
With diverse quantum hardware including superconducting qubits, trapped ions, neutral atoms, and photonic qubits, various quantum software packages focus on different aspects of quantum computing. This may include circuit synthesis or optimizing complex design processes including qubit allocation, auxiliary qubit reuse, error mitigation, or quantum error correction. These quantum software packages are developed in different programming models and have parallelism models supporting different backends. For example, CUDA-Q supports both task-based and distributed parallel circuit execution models and provides multi-GPU, multi-node state vector and tensor network simulation backends on Nvidia GPUsNVIDIA. (see Appendix G for details). Pennylane Lightning, as another example, provides state vector simulators that can be executed on both AMD and Nvidia GPUsAsadi et al. (2024).
@@
To enable seamless invocation of quantum kernels from different quantum SDKs within HPC applications developed in C/C++/Fortran, an API-CPE Quantum Interface Library (CPE-QIL) is implemented in C, which can be seamlessly interfaced with Fortran applications using the iso_c_binding module. Application developers can use high-level, portable invocation of quantum algorithm libraries from a variety of third-party SDKs within a classical HPC application in their programming language of choice. Figure 32 provides a schematic for a hybrid quantum–HPC development workflow within CPE with the quantum interface library.
@@
Circuit synthesis and execution time in quantum computing can vary significantly based on the complexity of the algorithm and the number of qubits involved. Different SDKs offer various levels of optimization and distinct approaches to handling quantum gates, scheduling, and resource allocation, all of which impact the time and efficiency of circuit synthesis and execution. Each quantum SDK (e.g., Qiskit and CUDA-Q) has its native gate set optimized for the underlying hardware. The choice of gate set impacts synthesis complexity since some SDKs might require multiple gates to synthesize specific operations efficiently, while others may directly support them. Decomposition of high-level operations into hardware-native gate sets is usually required. For example, synthesizing a rotation or controlled operation might require multiple CNOT gates and rotations, affecting both gate count and execution time. Many SDKs provide optimization levels to minimize circuit depth, gate count, or overall execution time. These optimizations directly influence both the synthesis time (by making trade-offs for synthesis overhead) and execution time on supported simulators and hardware. An example of time ranges is given in Section VI.1, where circuit synthesis for Trotterization is found to take on the order of milliseconds.
@@
Quantum interface library with API extensions: The quantum interface library extension within CPE has the following advantages: 1) the ability to support and interact with a range of quantum SDKs and backends from different vendors through a standardized interface; 2) data and functionality of other software systems are available while implementation details are abstracted, facilitating efficient and secure application development; 3) support for compiled languages commonly used in massively parallel HPC applications which could allow large-scale system modeling and computation with large datasets; and 4) direct utilization of the existing HPE Cray MPI, which offers "GPU aware" MPI support and heterogeneous workload and resource managers such as Slurm and PBS (Portable Batch System) Slurm. ; OpenPBS. .
@@
The CPE quantum interface library has been utilized to deliver two hybrid quantum–HPC applications in both Python and C/C++ with quantum kernels for circuit synthesis and execution provided by Classiq’s Python-based quantum SDK Classiq Technologies Ltd. . One example consists of solving linear systems of equations with the Harrow–Hassidim–Lloyd (HHL) algorithm Jose Morrell Jr et al. (2021). The solution was compared with the CPE BLAS library and showed less than 2% deviation. Another example considers the quantum approximate optimization algorithm (QAOA) for solving the MaxCut problem of partitioning a large graph into smaller sub-graphs, which can each be represented on current quantum devices. The workflow is well suited for a hybrid classical–quantum execution on supercomputers, where various sub-problems can also be solved classically if there is an advantage. Initial investigations have been conducted by HPE and Classiq and published at IPDPS24 Esposito and Danzig (2024). For a simplified problem, this hybrid workload was demonstrated at ISC24 using a 20-qubit IQM quantum device accessed from the LUMI supercomputer in Finland. Understanding the latency implications in accessing a remote quantum device was one of the main goals of this investigation; the communication overhead was typically found to be on the order of seconds. This latency could be reduced for tightly integrated machines.
@@
Figure 33 presents a detailed illustration of hybrid quantum–HPC application development and execution within CPE with the quantum interface library and HPC workload management. Within CPE, users can automatically utilize debugging and profiling tools as well as math and communication libraries with chosen compilers. These compilers (for Fortran, C, and C++) are designed to extract maximum performance from a variety of architectures like ARM and x86-64 and devices like AMD and Nvidia GPUs. In addition, users have access to the HPE Cray Message Passing Interface (MPI), a highly scalable implementation for collective communications. Applications developed in C/C++/Fortran are compiled and linked within CPE, potentially including other libraries such as the Cray Science and Math Libraries (CSML) if the HPC application requires it. The quantum interface library is linked as a shared library during the application build process. Hybrid executables can be submitted and scheduled for parallel execution by a workload manager such as Slurm or PBS. At runtime, the interface library routes the quantum API calls from the application to the respective vendor-specific quantum SDKs with appropriate data handling (e.g., parameters and return values). When a quantum API call provided by the interface library is invoked by the application at runtime:
@@
  • The quantum circuit is executed (either on a quantum device or simulator) and results are converted and passed back to the application as return values of the quantum API.
@@
Figure 33: Development and execution of hybrid quantum–HPC applications in C/C++/Fortran through the Quantum Interface Library within the HPE Cray Programming Environment (CPE). A hybrid executable is dispatched to a workload manager (e.g. Slurm or PBS) for parallel execution. Remote visualization through web applications can be enabled with a different protocol and configurations for secure connection to clusters.
@@
Quantum compiler and runtime extensions: With increasing qubit counts, higher gate fidelity, and improved coherence times and scalability, compilation bottlenecks and latency between classical and quantum components become more apparent. When scaling to large circuit sizes with $\sim$100 or more qubits, declarative frameworks struggle to handle compilation bottlenecks and latency between classical and quantum components in the program, even for NISQ systems. Leveraging classical compilation tool chains such as Clang/LLVM is crucial for developing large-scale hybrid quantum–HPC workloads, and research efforts are moving in this direction Litteken et al. (2020); Chong et al. (2017).
@@
To enable performant language-level support for quantum constructs and to address bottlenecks in compile-time with increasing circuit size, we have on-going efforts to build extensions on top of classical compilation tool chains. Given the diversity of pulse-level quantum instructions by different quantum hardware vendors, the quantum assembly language OpenQASM and LLVM IR with its extension to Quantum Intermediate Representation (QIR) Microsoft. are adopted as a middle ground for comparability with different quantum hardware and software. We leverage codegen modules to emit machine-specific native code for target architectures based on the LLVM IR/QIR produced by different quantum software front ends such as CUDA-Q. CUDA-Q provides the NVQ++ compiler for quantum kernels lowering to QIR eventually, as well as a standard library of quantum algorithmic primitives with upcoming support for quantum error correction primitives. Generated code, when linked with appropriate quantum vendor-provided runtime libraries, can be executed on target quantum devices or simulator backends in a quantum-backend-retargetable manner (see Appendix G for details).
@@
Dynamic code generation techniques are adopted to generate code at runtime based on specific characteristics of the target quantum devices. In doing so, the compiler can optimize the execution of quantum programs for different hardware generations or architectures. This low-level integration at the IR level will ensure hardware support, software compatibility, and optimal circuit compilation and execution performance for large-scale quantum–HPC workload development.
@@
Multi-QPU workload distribution hypervisor: The final extension to the HPC programming environment aims to tackle the problem of quantum workload distribution. For quantum computing to operate at scale, it will be necessary to efficiently parallelize over many QPUs, possibly of different hardware types, integrating them as coprocessors within an HPC framework. We enable this integration by developing a novel adaptive circuit knitting strategy serving as a hypervisor for classical and quantum communication between distributed classical and quantum compute nodes. Unlike traditional circuit knitting, this strategy uses machine learning at multiple levels to learn a quantum circuit decomposition capturing maximal quantum entanglement while enabling high-performance communication at scale. By integrating scalable message passing techniques within an adaptive adaptive circuit knitting approach, we could efficiently learn how to perform distributed quantum simulation or distributed quantum machine learning over quantum data generated by quantum processors themselves. We introduce this approach in the next section.
V.2 A programming framework for heterogeneous quantum–classical supercomputing
A major challenge in heterogeneous quantum–-classical computing is the development of software platforms that enable efficient programming of quantum–HPC systems. Many near- and mid-term algorithms and workflows require the simultaneous and efficient utilization of CPUs, GPUs, and QPUs. Achieving this level of performance requires accurate resource estimation and the development of schedulers capable of balancing algorithms and workloads across accelerators to maximize overall system utilization. One major obstacle is the access to hardware specifications and software libraries developed by quantum hardware vendors, which complicates integration. Additionally, the lack of standardized interfaces and protocols further hinders the seamless integration of QPUs into existing HPC systems.
Although various quantum–HPC platforms could be developed, those designed for near-term critical applications (e.g., quantum error correction) must be built to inherently support low-latency execution and tight coordination between GPUs and QPUs. A key concept for such software platforms is the use of kernels as units of work that can be executed on either GPUs or QPUs and efficiently transferred between them. CUDA-Q is a particularly promising software solution due to its architecture, which natively supports GPU-based acceleration and kernel execution. This makes it well-suited for a tightly coupled heterogeneous quantum–GPU system, like DGX Quantum, that requires efficient resource allocation and load balancing across quantum and classical resources.